OpenStudy (anonymous):

I need some help guys !! Solve using substitution: 5y – x – 6 = 0 -3x + 6y = -9

OpenStudy (anonymous):

Put one of those equations in terms of x, so rearrange terms so that x appears only on the left-hand side, then substitute this x-equation into the x variable in the second equation. Then you can solve for y. Once you know y, then solve for x.

OpenStudy (anonymous):

how do we do that

OpenStudy (anonymous):

firstly make one of the variable the subject say x in 5y-x-6=0 x=5y-6 then subtitute5y-6 in the 2nd equation -3(5y-6)+6y=-9 that makes -15y+18+6y=-9 then rearagnge to solve for y

OpenStudy (anonymous):

then u can solve for x using y you have just solved for.

OpenStudy (anonymous):

okay thanks